# Making Language Model a Hierarchical Classifier and Generator

### v1

@@ 1 Introduction
Figure 1: Paradigm of HdLM compared to a conventional decoder-only language model. In contrast to the last layer decoding, HdLM employs different layers to decode different levels of texts.
In this paper, we propose a new type of decoder called Hierarchical decoding Language model (HdLM), combining the model-wise and layer-wise prototypes. We inherit the strong semantic understanding and generation capabilities of modern LLM, but allow its multiple layers to decode plausible textual contents. An earlier layer may produce more prerequisite, coarse-grained expressions which also work for the context of resulted, fine-grained expressions, which are decoded by later layers (Figure 1). Such type a language model can automatically produce reasonable performance on sequential pipelines of classification and generation tasks, which are verified by our substantial experiments. Comparing with the conventional reasoning LLM, HdLM also enjoys computational savings for both training and inference, which are both verified by theoretical derivation and empirical observations. Our code and model can be found on https://github.com/coswindywang/HdLM. Our main contributions can be summarized as follows:
(1) We propose HdLM, which can deal with different types of hierarchical tasks, including hierarchical text classification, classification-guided generation, and hierarchical text generation.
(2) We provide a theoretical analysis on the computational benefit HdLM, and also a discussion on its convergence.
(3) We conduct substantial experiments and corresponding analysis to verify HdLM’s performance, on both classification and generation metrics, as well as its out-of-domain generalization capability.
@@ 2.1 Problem Formulation
where $\mathcal{T}$ denotes a uni-step subtask. Here we further argue $\mathcal{T}$ can be generally classified into two categories: the classification task $\mathcal{C}$ and the generation task $\mathcal{G}$. As indicated in Figure 2, the following different paradigms can be summarized from hierarchical textual tasks:
@@
Classification-guided generation (CgG):
a classification result is first required, then the final response is generated based on that classified label. This paradigm can be denoted as $\mathcal{C}\rightarrow\mathcal{G}$.
Hierarchical text generation (HTG):
the final response is posterior inferenced by intermediate generations, such as the chain of thought, reasoning, or thoughts. This paradigm is denoted as $\mathcal{G}\rightarrow\mathcal{G}$.
Other hierarchical tasks:
there might be even more completed hierarchical tasks. For example, the generative classification ($\mathcal{G}\rightarrow\mathcal{C}$) and ‘think, classify and act’ ($\mathcal{G}\rightarrow\mathcal{C}\rightarrow\mathcal{G}$). We leave these types of tasks for future works.
@@
Figure 2: Typical paradigms (hierarchical text classification, classification-guided generation, hierarchical text generation) of hierarchical text tasks. $\mathcal{C}$ denotes classification and $\mathcal{G}$ denotes generation.
@@ 2.2 Architecture
In this section, we propose an adjusted architecture based on the decoder-only transformer, which can also deal with hierarchical textual tasks by post-hoc adaptations. We first make a reasonable assumption that the depth of hierarchical textual task is smaller, i.e., $D<K$. Then we select $D-1$ intermediate layers with their indices satisfying:
@@
These layers, along with the final layer, are used to decode $D$ responses. To achieve this objective, we replicate the language heads in the $K$-th layer to the $D-1$ intermediate layers, with the parameter randomly initialized.
@@ 2.3 Training
The $D-1$ intermediate layers can not decode reasonably with newly added heads. To educate them to decode proficient language, post-hoc adaptation is needed.
@@
with the finetuning loss $\mathcal{L}_{d}$ implemented on each decoding layer. The final loss can then be the linear combination of them: $\mathcal{L}=\sum_{d=1}^{D-1}f_{d}\mathcal{L}_{d}+\mathcal{L}_{D}$ with the list of $\{f_{1},\cdots,f_{D-1},1\}$ as the loss weights. During the training, we implement the causal mask for both prior tokens and subsequent responses. Figure 3 exhibits the detailed masking matrices for all decoding layers.
Figure 3: The pipeline of HdLM, with exampled $D=3$ and $K=5$. (a): A depth=3 hierarchical task on a conventional LM. (b): Replicate the language heads of the $K$-th layer to layer $k_{1},k_{2}$. (c): Training HdLM with the fine-tuning loss of $r_{d}$ on the $k_{d}$-th layer. (d): Inference HdLM on the $k_{d}$-th layer sequentially with decoded $\hat{r}_{d}$ and corresponding length $\hat{L}_{d}$.
@@ 2.4 Inference
3 Theoretical analysis
This section theoretically analyzes the computation savings of HdLM compared to a conventional LLM. We then briefly discuss the training convergence of HdLM with the connection between the index of the first decoding layer and the task depth.
3.1 Computational complexity analysis
In this subsection, we compare the computation FLOPs of our HdLM with a standard LLM, and verify that HdLM results in both training and inference reduction. To make the formulation simple and clear, here we discuss the 2-depth hierarchical task, i.e., $D=2$.
Parameter definitions.
The input sequence length $L$, the output sequence lengths $L_{1}$ and $L_{2}$, the hidden dimension $E$, indexes of two decoding layers $k_{1}<K$ and $k_{2}=K$, and the expand ratio of FFN dimension is $c$. We use $\mathcal{F}$ to denote the FLOPs.
Training FLOPs savings.
We begin with the forward FLOPs of a standard transformer from Equation 20 derived in Appendix: $\mathcal{F}_{\text{train}}\rightarrow 3(8+4c)E^{2}KL$  [^1] Here we omit the batch size term $B$ since different samples’ computation within a batch is parallel.. The LLM baseline simply treats the input, level-1 output and level-2 output as a flat sequence, with all layers engaged:
$\mathcal{F}_{\text{train}}^{\text{baseline}}\approx 3fK(L+L_{1}+L_{2})$ (8)
where we use $f$ to represent the fixed coefficient $(8+4c)E^{2}$ for simplicity. One the other hand, the training FLOPs of HdLM can be expressed as
$\mathcal{F}_{\text{train}}^{\text{{HdLM}}}\approx 3fk_{1}(L+L_{1})+3f(K-k_{1})(L+L_{1}+L_{2})$ (9)
The training FLOPs savings then becomes
$\mathcal{F}_{\text{train}}^{\text{baseline}}-\mathcal{F}_{\text{train}}^{\text{{HdLM}}}\approx 3fk_{1}L_{2}>0$ (10)
Inference FLOPs savings.
The inference FLops of LLM baseline can be expressed as
$\mathcal{F}_{\text{infer}}^{\text{baseline}}=fK\sum_{j=1}^{L_{1}+L_{2}}(L+j-1)$ (11)
from Equation 21 derived in Appendix. While for HdLM, the inference FLOPs becomes
$\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}=fk_{1}\sum_{j=1}^{L_{1}}(L+j-1)+f(K-k_{1})\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)$ (12)
and it is easy to further prove $\mathcal{F}_{\text{infer}}^{\text{baseline}}-\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}>0$ (detailed derivation in Appendix).
3.2 Convergence of decoder with arbitrary layers
This subsection briefly discusses the convergence issue, i.e., how can we ensure any implementation decoding layer (with the index of $k_{d}$), has enough depth to learn the target semantic hierarchy (with depth $d$)? Our derivation is mainly motivated by the Minimum number of layers Theorem proposed in [11]
Theorem 1.
The last position in the sequence cannot retrieve the embedding vector $x_{D}$ of the target token with $K$ transformer layers if $K<\log_{3}(2D)$.
Due to the page limit, here we leave the detailed assumption in [11] and the connection between retrieval and classification in the Appendix. Based on its conclusion, here we prove our Corollary 2:
Corollary 2.
Assume Theorem 1 holds in $\mathcal{C}$ or $\mathcal{G}$, then for any $d\leq D$, the first $k_{d}$ layers of transformer can always learn the top $d$-depth task information.
Proof.
Given the strict increasing, non-repeat integer sequence $\{k_{d}\}$ with $d\in\{1,2,\cdots,D\},k\in\{1,2,\cdots,K\},D\leq K$, it is obvious to have $k_{d}\geq d$. Since $3^{d}>2d$ holds for $d>=1$, then we have $k_{d}\geq d>\log_{3}{2d}$, which contradicts the situation in Theorem 1. ∎
Corollary 2 ensures the theoretical feasibility of our approach. In the next section, we also verify this conclusion from empirical results.
@@ 4 Experiment
In this section, we consider answering the following research questions: RQ1: If the training of HdLM can converge with an arbitrary setting of $k$? RQ2: Can HdLM perform well on different types of hierarchical textual tasks and outperform conventional baselines? RQ3: Can HdLM generalized well on out-of-domain hierarchical textual datasets? RQ4: How to determine the key parameters and different components of HdLM? RQ5: What is the connection of HdLM with previous standardized or test-time scaling baselines, such as reasoning, multi-hop, or cascade solutions?
To address these issues, we first list the datasets we use for different types of hierarchical tasks, then the experimental settings and evalulation metrics, and followed by the main results. We finally provide thorough discussion on the open-domain performance, the ablation and sensitivity studies, and the actual computational savings.
@@ 4.1 Datasets
CgG ($\mathcal{C}\rightarrow\mathcal{G}$).
We employ the emotional support conversation (ESC) scenario, in which each turn of conversation is annotated with the user’s emotion and the assistant’s response strategy. The model needs to sequentially classify the user emotion or the response strategy to further enhance the response quality. We utilize ESconv [14] as both training and in-domain (ID) test sets. To further validate the response generation, EmpatheticDialogues [15] is employed as an out-of-domain (OOD) test set which lacks of the strategy annotation.
HTG ($\mathcal{G}\rightarrow\mathcal{G}$).
Here we examine two typical famous theory-of-mind (ToM) tests, the Sally-Anne false belief experiment, and the Vignette-based Problem. These cognitive tests contain misleading or disturbing information and the agent needs to reason from the ego-centric perspective. In these scenarios, the response is the (thought, answer) pair with depth $D=2$. We train HdLM on the ToMI [16] and BigToM [17] benchmarks, while zero-shot evaluate the model on the Vignette-based test provided by [18].
@@ 4.2 Setting
F1-related scores including Micro-F1 and Macro-F1 are used. Micro-F1 considers the overall precision and recall of all instances, while Macro-F1 equals the average F1-score of labels. For CgG tasks, we also provide the classification accuracy, and the preference $bias$ as defined by [21] based on Bradley-Terry model [22]  [^2] Detailed formula in the Appendix. Smaller $bias$ means better.. For HTC tasks, we provide the metric results on the bottom level of labels.
Generation metrics.
We utilize the famous metrics of BLEU-2 (B-2), Rouge-L (R-L) and CIDEr. BLEU-2 (B-2)[23] first compute the geometric average of the modified $n$-gram precisions, then calculate the BP and BLEU-2 result. Rouge-L (R-L)[24] uses LCS-based F-measure to estimate the similarity between two summaries. CIDEr [25] calculates the cosine similarity from the average of different $n$-grams.
@@ 4.3 Results
Table 1 shows the F1 results of WoS (depth=2) and DBpedia (depth=3), compared to previous state-of-the-art HTC baselines. Our HdLM performs the best on both WoS and DBpedia, indicating it has a reasonable semantic hierarchy comprehension and conducts accurate bottom-level classification. For typical cases of HdLM on HTC, see Appendix.
@@
HdLM (ours) 88.40 87.54 96.73 96.37
Classification-guided Generation.
For ESC sc, we consider baselines including Direct (direct inference), Refine (revise the initial response immediately), Self-Refine [32] (multi-hop inference with refinement on the first-time generation), CoT [33] and the finite state machine (FSM) [34].
Table 2 shows that HdLM also has a reasonable performance on both classification and generation metrics, either the largest or the second-largest. HdLM is also robust to different classification paradigms (with or without emotion) and generalizes well on OOD situations (EmpatheticDialogues). For typical cases of HdLM on CgG, see Appendix.
Table 2: Results of classification-guided generation on ESconv (IID) and EmpatheticDialogues (OOD), including classification metrics such as Accuracy (ACC), Macro-F1 (MaF1) and $bias$, and generation metrics such as BLEU-2 (B-2), ROUGE-L (R-L) and CIDEr. Among the task subscripts, ‘emo’ denotes emotion, ‘stra’ denotes strategy, and ‘resp’ denotes response. The best results of each model are bolded and the second best are underlined.
Method ESconv ($\mathcal{C}_{emo}\rightarrow\mathcal{C}_{stra}\rightarrow\mathcal{G}_{resp}$) EmpatheticDialogues ($\mathcal{C}_{stra}\rightarrow\mathcal{G}_{resp}$)
ACC $\uparrow$ MaF1 $\uparrow$ $bias$downarrow B-2 $\uparrow$ R-L $\uparrow$ B-2 R-L CIDEr
Direct 11.80 10.26 1.61 3.47 10.64 3.09 9.91 1.60
+ Refine 17.08 11.07 1.27 3.10 6.13 2.56 9.12 0.42
+ Self-Refine 17.58 13.61 1.92 3.34 9.71 3.08 9.91 1.56
+ CoT 15.32 10.38 1.69 3.16 10.50 2.91 9.79 1.37
+ FSM 17.37 11.15 0.81 4.12 11.83 3.33 10.80 2.96
SFT 30.60 21.29 1.28 6.97 16.59 4.10 10.88 6.84
+ CoT 30.80 17.70 1.35 6.51 15.00 5.05 14.06 12.01
+ FSM 28.83 18.36 1.32 7.57 17.42 4.69 14.28 10.12
+ HdLM (ours) 33.41 21.65 0.89 7.54 18.13 5.48 14.53 11.95
Hierarchical Textual Generation.
In this experiment, we compare with different reasoning LLMs (CoT[33], SimToM [35], Quiet-STaR [6]), and different base models (Llama3-8B-Instruct, Mistral-7B-Instrct, GPT3 and GPT4). To collect the finetuning dataset, we inference SimTom [35] on ToMI and BigToM and employ its reasoning contents as our $r_{1}$. Therefore, our HdLM can also be considered as a distilled reasoning LLM.
Table 3: Results of Sally-Anne false belief experiments (ID) and vignetted-based tests (OOD). Values of results are in percentage.
Base Method ToMI BigToM Vignette Scores
models with much larger size:
GPT-3 direct - - 37.5
GPT-4 direct 92.5 66.5 46.9
GPT-4 CoT[33] 95.5 74.4 -
GPT-4 SimTom [35] 95.0 87.8 -
models with similar size:
Mistral-7B-Instruct direct - - 40.2
Mistral-7B-Instruct Quiet-STaR [6] - - 11.1
Llama3-8B-Instruct direct 22.2 71.3 23.8
Llama3-8B-Instruct SFT 43.2 77.7 -
Llama3-8B-Instruct HdLM (ours) 98.2 99.4 48.3
[…]
Table 3 indicates that HdLM has the highest scores, surpassing prompt and finetuned baselines. HdLM even outperforms GPT-3 and GPT-4 [1] which has a much larger size. Also, Quiet-STaR (based Mistral-7B-instruct) although also has an internal thinking mechanism, fails to capture the Vignette-based scenario, since it is more focused on math reasoning. For typical cases of HdLM on HTG, see Appendix.
4.4 Discussion
To further address the RQs, here we provide a series of extra studies. To make the discussion clear, here we simplify the hierarchical scenario by constraining its depth $D=2$. Subsequently, the responses are $(r_{1},r_{2})$ and the decoding layer indexes are $(k,K)$ where we simplify $k$ for $k_{1}$.
@@ 4.4 Discussion
Table 4 shows the ablation result on WoS. HdLM still holds the best results which indicates our framework is rigorous. Furthermore, we also compare HdLM to two imaginary solutions
@@
both of which are apparently not applicable in the practice. Even so, HdLM’s performance is still close to them, indicating a high-level comprehension of label hierarchy.
Ablation on hierarchical levels.
HdLM trained by ESconv has the paradigm $\mathcal{C}_{emotion}\rightarrow\mathcal{C}_{strategy}\rightarrow\mathcal{G}_{response}$, which suggests the final response is grounded by its decoded emotion and strategy results. To illustrate this sequential reasoning process effective, we compare to two ablations, w/o emotion and w/o strategy, in Table 5. HdLM performs the best almost on all metrics.
@@
w/ head cascade multi-hop sft@$K$ HdLM true-1 true-1 & true-2 candidates
@@
Table 5: Ablation results of classification-guided generation on ESconv (ID) and EmpatheticDialogues (OOD), including classification metrics (on strategy) such as Accuracy (ACC), micro-F1 (miF1) and $bias$, and generation metrics such as BLEU-2 (B-2), ROUGE-L (R-L) and CIDEr.
Method ESconv (ID) EmpatheticDialogues (OOD)
ACC $\uparrow$ miF1 $\uparrow$ $bias$downarrow B-2 $\uparrow$ R-L $\uparrow$ B-2 $\uparrow$ R-L $\uparrow$ CIDEr $\uparrow$
w/o emotion 30.34 21.14 0.43 6.83 16.35 5.32 14.43 11.92
w/o strategy N/A N/A N/A 7.37 17.3 5.47 15.60 7.74
 HdLM 33.41 21.65 0.89 7.54 18.13 5.48 14.53 11.95
Selection of the decoding layer.
Figure 4 compares the metrics on WoS and ESconv, with respect to different choices of $k$. When $k$ is small, the performance is relatively low since the layer-depth of HdLM is not enough for $r_{1}$ understanding and generation. As $k$ increases, performances also improve. However, when $k$ is close to $K$, they degrade again the layer-depth for $r_{D}$ starts to be limited. From these results, we choose $k=25$ for WoS and $k=28$ for ESconv.
Figure 4: Performance sensitivities on WoS (left) and ESconv (right) with different decoding layer selection ($k$).
@@
We also calculate the computational times in the experimental environment. For different values of $L_{2}$ (128, 256, 512), the computation reduction of HdLM grows linearly (7.03%, 18.03%, 26.25%), no matter what the input length is. On the other hand, for different $k$ (16, 20, 24), the computation reduction of HdLM also grows linearly (7.43%, 12.09%, 17.85%). These observations verify the theoretical conclusion. Detailed plots and corresponding error bars are included in the Appendix.
Other discussions.
Further analysis are included in the Appendix D.5 due to page limit. Such studies include loss weight sensitivity and loss curves. Summarizing from these results, we suggest that the generation subtask may need larger layer depth and higher loss weights to converge, compared to the classification subtask. HdLM also generalizes reasonably to open-domain planning cases.

### v2

@@ 1 Introduction
In this paper, we propose a new type of decoder called Hierarchical Decoding Classifier (HDC), combining the model-wise and layer-wise prototypes. We inherit the strong semantic understanding and generation capabilities of modern LLM, but allow its multiple layers to decode plausible textual contents. An earlier layer may produce more prerequisite, coarse-grained expressions which also work for the context of resulted, fine-grained expressions, which are decoded by later layers. Such type a language model can automatically produce reasonable performance on sequential pipelines of classification and generation tasks, which are verified by our substantial experiments. Comparing with the conventional reasoning LLM, HDC also enjoys computational savings for both training and inference, which are both verified by theoretical derivation and empirical observations. Our main contributions can be summarized as follows:
(1) We propose HDC, which can deal with tasks of hierarchical text classification.
(2) We provide a theoretical analysis on the computational benefit HDC, and also a discussion on its convergence.
(3) We conduct substantial experiments and corresponding analysis to verify HDC’s performance, as well as its out-of-domain generalization capability.
@@ 2.1 Problem Formulation
where $\mathcal{T}$ denotes a uni-step subtask. Here we further argue $\mathcal{T}$ can be generally classified into two categories: the classification task $\mathcal{C}$ and the generation task $\mathcal{G}$. As indicated in Figure LABEL:fig:task, the following different paradigms can be summarized from hierarchical textual tasks:
@@ 2.2 Architecture
In this section, we propose an adjusted architecture based on the decoder-only transformer, which can also deal with hierarchical textual tasks by post-hoc adaptations. We first make a reasonable assumption that the depth of hierarchical textual task is smaller, i.e., $D<K$. Then we select D1D-1 intermediate layers with their indices satisfying:
@@
These layers, along with the final layer, are used to decode $D$ responses. To achieve this objective, we replicate the language heads in the $K$-th layer to the D1D-1 intermediate layers, with the parameter randomly initialized.
@@ 2.3 Training
The D1D-1 intermediate layers can not decode reasonably with newly added heads. To educate them to decode proficient language, post-hoc adaptation is needed.
@@
with the finetuning loss $\mathcal{L}_{d}$ implemented on each decoding layer. The final loss can then be the linear combination of them: $\mathcal{L}=\sum_{d=1}^{D-1}f_{d}\mathcal{L}_{d}+\mathcal{L}_{D}$ with the list of $\{f_{1},\cdots,f_{D-1},1\}$ as the loss weights. During the training, we implement the causal mask for both prior tokens and subsequent responses.
@@ 3.2 Setting
F1-related scores including Micro-F1 and Macro-F1 are used. Micro-F1 considers the overall precision and recall of all instances, while Macro-F1 equals the average F1-score of labels. For CgG tasks, we also provide the classification accuracy, and the preference biasbias as defined by [21] based on Bradley-Terry model [22]  [^1] Detailed formula in the Appendix. Smaller biasbias means better.. For HTC tasks, we provide the metric results on the bottom level of labels.
@@ 3.3 Results
Table 1 shows the F1 results of WoS (depth=2) and DBpedia (depth=3), compared to previous state-of-the-art HTC baselines. Our HDC performs the best on both WoS and DBpedia, indicating it has a reasonable semantic hierarchy comprehension and conducts accurate bottom-level classification. For typical cases of HDC on HTC, see Appendix.
@@
HDC (ours) 88.40 87.54 96.73 96.37
@@
Table 2 shows the ablation result on WoS. HDC still holds the best results which indicates our framework is rigorous. Furthermore, we also compare HDC to two imaginary solutions
@@
both of which are apparently not applicable in the practice. Even so, HDC’s performance is still close to them, indicating a high-level comprehension of label hierarchy.
@@
w/ head cascade multi-hop sft@$K$ HDC true-1 true-1 & true-2 candidates
@@
We also calculate the computational times in the experimental environment. For different values of $L_{2}$ (128, 256, 512), the computation reduction of HDC grows linearly (7.03%, 18.03%, 26.25%), no matter what the input length is. On the other hand, for different $k$ (16, 20, 24), the computation reduction of HDC also grows linearly (7.43%, 12.09%, 17.85%). These observations verify the theoretical conclusion. Detailed plots and corresponding error bars are included in the Appendix.